OpenStudy (anonymous):

what is 2/5 of 15 and how do I show work?

OpenStudy (texaschic101):

2/5 of 15 --- " of " means multiply 2/5 * 15 = 30/5 = 6 6 = 2/5(15) 6 = 30/5 6 = 6 (correct)

OpenStudy (anonymous):

thank you so much @texaschic101

OpenStudy (texaschic101):

sure thing :)